OpenStudy (anonymous):

The sum of two numbers is 55 . One number is 4 times as large as the other. What are the numbers?

rishavraj (rishavraj):

see let no be "x" so other no = 4x x + 4x = 55 solve it

sammixboo (sammixboo):

One number = x ????? = 55 4x + x = 55 Since one is 4 times larger, you must multiply that number (x) by 4

OpenStudy (anonymous):

is it 13.75?

rishavraj (rishavraj):

hmm how???? 4x + x = 55 5x = 55 x = ?????

OpenStudy (anonymous):

11

sammixboo (sammixboo):

Check. 4x + x = 54 13.75*4 + 13.75 = 55 55 + 13.75 = 55 68.75 =/= 55

rishavraj (rishavraj):

yup one no i.e x = 11 and other 4x = 4*11 = 44

sammixboo (sammixboo):

Check 4x + x = 55 4*11 + 11= 55 44 + 11 = 55 55 = 55 Right!